OpenStudy (anonymous):

adam can paint a room in 5 hours by himself. Audrey can do the same job in 6 hours. If they work together how long will it take

OpenStudy (anonymous):

(x/5)+(x/6)=1

OpenStudy (anonymous):

Can you solve from here?

OpenStudy (anonymous):

no i cant

OpenStudy (anonymous):

I have go restroom,I be back

OpenStudy (anonymous):

Denominator is: 30 6x+5x=30 11x =30 x= 30/11 x= 2.73 hours 2 hours and 44 minutes